OpenStudy (anonymous):

what is the square root of 100v^3

OpenStudy (jdoe0001):

what's \(\sqrt{100}\)?

OpenStudy (anonymous):

10v^2sqrt(v)

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

because 10 * 10 = 100 and when you sqrt(v^3) you get v^2

OpenStudy (anonymous):

and get left out with v

OpenStudy (anonymous):

so it becomes 10v^2sqrt(v)

OpenStudy (jdoe0001):

\(\bf \sqrt{v^3} \implies \sqrt{v^2v} \implies v\sqrt{v}\)

OpenStudy (anonymous):

ohh yeah nvm

OpenStudy (anonymous):

just follow jdoe0001

OpenStudy (anonymous):

i not not doing well today

OpenStudy (jdoe0001):

"Don't let life discourage you; everyone who got where he is had to begin where he was." Richard L. Evans

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[what is the \sqrt[3]{512k^2}\]